Two point charges are separated by a distance r. If the separation is reduced by a factor of 1.5, by what factor does the electric force between them change? 


1
Expert's answer
2022-01-31T11:27:28-0500

Now We know that

Electric force

"F=\frac{kq_1q_2}{r^2}"

"F\alpha\frac{1}{r^2}"

Another case

"F'\alpha\frac{1}{r'^2}"

Both equation We can written as

"\frac{F}{F'}=\frac{r'^2}{r^2}"

"r'=\frac{r}{1.5}"

"\frac{F}{F'}=\frac{1}{1.5^2}"

"F'=2.25F"